How do I do x+y=34 and -2x+4y=110 by elimination?
Answer by scott8148(6628)

one of the variables has to have the same coefficient in both equations
___ so that adding or subtracting will cancel (eliminate) the variable

multiplying the 1st equation by 2 gives ___ 2x + 2y = 68

adding this to the 2nd equation cancels the x-terms ___ 2x + 2y - 2x + 4y = 68 + 110 ___ 6y = 178

solve for y, then substitute back to find x